# LinguaLIFT: An Effective Two-stage Instruction Tuning Framework for Low-Resource Language Tasks

### v1

@@ Abstract
Large language models (LLMs) have demonstrated impressive multilingual understanding and reasoning capabilities, driven by extensive pre-training multilingual corpora and fine-tuning instruction data. However, a performance gap persists between high-resource and low-resource language tasks due to language imbalance in the pre-training corpus, even using more low-resource data during fine-tuning. To alleviate this issue, we propose LinguaLIFT, a two-stage instruction tuning framework for advancing low-resource language tasks. An additional language alignment layer is first integrated into the LLM to adapt a pre-trained multilingual encoder, thereby enhancing multilingual alignment through code-switched fine-tuning. The second stage fine-tunes LLM with English-only instruction data while freezing the language alignment layer, allowing LLM to transfer task-specific capabilities from English to low-resource language tasks. Additionally, we introduce the Multilingual Math World Problem (MMWP) benchmark, which spans 21 low-resource, 17 medium-resource, and 10 high-resource languages, enabling comprehensive evaluation of multilingual reasoning. Experimental results show that LinguaLIFT outperforms several competitive baselines across MMWP and other widely used benchmarks.  [^1] Our code and data will be released once accepted.
LinguaLIFT: An Effective Two-stage Instruction Tuning Framework for Low-Resource Language Tasks
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated impressive understanding and reasoning capabilities, opening up a new paradigm for artificial intelligence (Wei et al., 2022; Zhao et al., 2023). These capabilities arise from self-supervised pre-training on large-scale multilingual data, followed by supervised fine-tuning on multilingual instruction data (Ouyang et al., 2022; Muennighoff et al., 2023; Chung et al., 2024). As a result, LLMs have achieved remarkable performance across various natural language processing tasks (Chen et al., 2024a, b; Zhang et al., 2024b, a), especially high-resource language mathematical reasoning tasks (Shi et al., 2023; She et al., 2024) and natural language inference tasks (Üstün et al., 2024; Lu et al., 2024).
@@
However, a substantial performance gap still exists between high-resource and low-resource language tasks (Zhu et al., 2024c), primarily due to the language imbalance within the large-scale multilingual corpora during pre-training, where high-resource languages dominate (Touvron et al., 2023). This gap remains unaddressed, even using more low-resource language data during the following fine-tuning Singh et al. (2024). As illustrated in Figure  1, while LLMs can generate the correct answer to the query in high-resource languages like English, they struggle with the same query in low-resource languages. These discrepancies emphasize a fundamental limitation in the understanding and reasoning capabilities of LLMs for low-resource languages.
To alleviate this issue, we propose a two-stage instruction tuning framework (LinguaLIFT) for enhancing low-resource language tasks. An additional language alignment layer is initially integrated into the LLM to adapt a pre-trained multilingual encoder. LinguaLIFT then improves multilingual alignment via code-switched fine-tuning, using the code-switched translation data generated from unsupervised multilingual alignment lexicons The second stage fine-tunes LLMs with English-only instruction data while freezing the language alignment layer, which enables LLMs to transfer the learned task-specific capabilities from English to low-resource languages. Additionally, we introduce a new benchmark named Multilingual Math World Problem (MMWP), spanning 21 low-resource, 17 medium-resource, and 10 high-resource languages to evaluate multilingual reasoning tasks comprehensively. Experiments demonstrate that LinguaLIFT significantly outperforms several strong competitive methods on the MMWP and other widely used benchmarks, such as MGSM (Shi et al., 2023), MSVAMP (Chen et al., 2023), XNLI (Conneau et al., 2018) and X-CSQA (Lin et al., 2021).
@@ 2 Related Work
Leveraging Powerful Cross-lingual Transfer Capabilities of Multilingual Models.
Recent studies have shown that multilingual models possess strong cross-lingual transfer capabilities (Conneau et al., 2020; Xue et al., 2021; FitzGerald et al., 2023; Chirkova and Nikoulina, 2024; Shaham et al., 2024; Chen et al., 2024c), benefiting the low-resource language community by enabling task transfer across multiple languages after fine-tuning on high-resource languages (Kew et al., 2023; Pfeiffer et al., 2020; Reimers and Gurevych, 2020; Pan et al., 2021; Feng et al., 2022).
Unlike previous approaches focusing on multilingual instruction tuning (Singh et al., 2024; Üstün et al., 2024; Köpf et al., 2024; Zhu et al., 2024b, a; Li et al., 2023a; Wei et al., 2023; Ranaldi et al., 2023), LinguaLIFT incorporates a language alignment layer into LLMs to adapt a pre-trained multilingual encoder through code-switched fine-tuning. While freezing the language alignment layer and a pre-trained multilingual encoder, the LLM is fine-tuned with English-only instruction data to transfer task-specific capabilities from English to low-resource language tasks.
@@
Recent efforts to improve multilingual mathematical reasoning for LLMs can be categorized into three ways: 1) Prompting close-source LLMs: Qin et al. (2023), Shi et al. (2023), and Huang et al. (2023) designed prompts for closed-source LLMs like ChatGPT, translating non-English contexts into English for reasoning. However, this approach is limited by translation quality and does not improve multilingual understanding or work well for open-source LLMs (Zhu et al., 2024b). 2) Instruction-tuning open-source LLMs: Chen et al. (2023), Chai et al. (2024), and Lai and Nissim (2024) adopted a translate-training method, translating English reasoning datasets into non-English and instruction-tuning LLMs. Zhu et al. (2024b) and She et al. (2024) proposed approaches to transfer mathematical reasoning capabilities from English to non-English. While these methods improve multilingual reasoning, they incur high translation costs and errors, making them impractical for low-resource languages. 3) Bridging existing skilled LLMs to multilingualism: Yoon et al. (2024) combines pre-trained multilingual models with skilled reasoning LLMs, but a performance gap persists between low-resource and high-resource languages. While Huang et al. (2024) shows significant improvement, it relies heavily on parallel and multilingual task data.
Different from existing studies, this paper proposes a novel two-stage instruction fine-tuning framework to enhance the reasoning capabilities of low-resource language tasks without relying on multilingual instruction data.
Figure 2: Overview of the proposed LinguaLIFT approach. Stage-I (Language-Align): A language alignment layer is introduced into the LLM to adapt the pre-trained multilingual encoder, thereby enhancing multilingual alignment through code-switched tuning. Stage-II (Task-Transfer): The LLM is fine-tuned on high-quality, English-only instruction data with the language alignment layer frozen, allowing LLM to transfer task-specific capabilities learned from English to low-resource languages.
@@ 3 Methodology
Figure 2 illustrates an overview of the proposed two-stage instruction tuning framework, LinguaLIFT. The core idea is to transfer task-specific capabilities learned from English to low-resource languages, leveraging the language alignment established by code-switched tuning. We then introduce the model architecture (§3.1), the language alignment stage (§3.2) and the task transfer stage (§3.3).
3.1 Model Architecture
Given a task-specific input $x$ with $l_{x}$ tokens, we use a multilingual pre-trained model to encode it into a moderately language-agnostic representation $X$, facilitating more effective cross-lingual understanding and transfer:
$X=Encoder(x),$ (1)
where $Encoder(\cdot)$ is the pre-trained multilingual encoder, and $X\in\mathbb{R}^{l_{x}\times d_{1}}$ is the final hidden state.
We introduce the language alignment layer, $Alignment(\cdot)$, to bridge the gap between the multilingual encoder and the LLM representation, implemented as a multi-layer perceptron (MLP). This layer maps $X$ to the input embedding space of the LLM:
$\hat{X}=Alignment(X),$ (2)
where $\hat{X}\in\mathbb{R}^{l_{x}\times d_{2}}$ is the alignment representation of $X$ on the input embedding space of the LLM.
To leverage the built-in capabilities of the LLM for instruction following, we embed the instruction context $q$ with $l_{q}$ tokens using the embedding layer $Embedding(\cdot)$:
$Q=Embedding(q),$ (3)
where $Q\in\mathbb{R}^{l_{q}\times d_{2}}$ presents the instruction context $q$ in the input embedding space of the LLM. Then, we concatenate the instruction context representation with the multilingual input representation:
@@ 3.1 Model Architecture
3.2 Language Alignment
In this stage, we enhance the model’s multilingual alignment through code-switched tuning. This involves two key parts: building multilingual alignment lexicons and enhancing alignment via code-switched translation instruction tuning.
Building Multilingual Alignment Lexicons.
Given the scarcity of data for low-resource languages, we adopt an unsupervised word translation method MUSE (Lample et al., 2018) to construct the multilingual alignment lexicons without parallel data. Specifically, we extract a set of unique words $\mathcal{W}=\{w\}_{i}^{N}$ from English monolingual corpus $\mathcal{D}$, where $N$ is the number of unique words, excluding named entities and stop words. We collect the translations for each word $w\in\mathcal{W}$ by retrieving the top nearest neighbor by projecting the word embeddings into the target space. Further implementation details are provided in Appendix A.1.
Enhancing Alignment with Code-Switched Translation Instruction Tuning.
We leverage the built multilingual alignment lexicons to generate code-switched inputs. Formally, this tuning aims to predict the English translation response tokens $y$ given the instruction context $q$ and code-switched input tokens $x_{l}$ in language $l$. The input embeddings for this task are represented as $(Q,\hat{X_{l}})$ from Equation (4). The optimization objective is outlined as follows:
[…]
where $T$ represents the sequence length of the target output $y$, $y_{<i}$ denotes the tokens preceding the $i$-th token, and $y_{i}$ is the $i$-th token of $y$. The trainable parameters include $\theta$, corresponding to the language alignment layer and the boundary tokens. $\sigma$ and $\phi$ represent the parameters of the multilingual encoder and LLM, respectively, which remain frozen during training to mitigate the risk of catastrophic forgetting. Finally, $\mathcal{L}$ denotes the set of target languages.
3.3 Task Transfer
After establishing language alignment in the first stage, we fine-tune LLM using high-quality, English-only task-specific instruction data while freezing the language alignment layer. This approach preserves the multilingual alignment and cross-lingual transferability capabilities acquired in the first stage, unlocking the potential of LLM to transfer strong capabilities learned from English to low-resource language tasks. Specifically, we fine-tune the parameters of the LLM to minimize the language modeling negative log-likelihood of the response $y$, given the combined instruction context and task-specific input $(Q,\hat{X}_{en})$:
$-\underset{\phi}{\arg\min}\sum_{i=1}^{T}\log p_{\tilde{\theta},\sigma,\phi}\left(y_{i}\mid(Q,\hat{X}_{en}),y_{<i}\right),$ (6)
where $\phi$ represents the trainable parameters of the LLM, and $\tilde{\theta}$ is initialized from the checkpoint of $\theta$ trained at the first stage. The multilingual encoder parameters $\sigma$ and $\tilde{\theta}$ are kept frozen throughout the training.
@@ 4 The MMWP Benchmark
We select a typologically diverse set of 48 languages, spanning 12 language families and 12 writing systems, with a range of resource levels—low, medium, and high. This includes 21 low-resource languages [^2] Afrikaans (af), Arabic (ar), Belarusian (be), Bengali (bn), Basque (eu), Gujarati (gu), Hausa (ha), Hindi (hi), Armenian (hy), Icelandic (is), Kannada (kn), Luxembourgish (lb), Macedonian (mk), Malayalam (ml), Marathi (mr), Nepali (ne), Slovak (sk), Swahili (sw), Tamil (ta), Telugu (te), Thai (th), 17 medium-resource languages [^3] Bulgarian (bg), Catalan (ca), Czech (cs), Danish (da), Finnish (fi), Croatian (hr), Hungarian (hu), Indonesian (id), Korean (ko), Norwegian Bokmål (nb), Polish (pl), Portuguese (pt), Romanian (ro), Slovenian (sl), Serbian (sr), Ukrainian (uk), Vietnamese (vi), and 10 high-resource languages [^4] German (de), English (en), Spanish (es), French (fr), Italian (it), Japanese (ja), Dutch (nl), Russian (ru), Swedish (sv), Chinese (zh). Additional selection details are provided in Appendix B.1.
Translation process.
To ensure the high quality of translations, we employed Google Translation System to translate the selected English questions into 47 other languages. Moreover, five annotators then post-edit and calibrate the translation, followed by quality estimation through both human and automatic evaluation, as detailed in Appendix B.2.
@@ 5.1 Datasets
We use the MMWP and latest multilingual benchmarks, MGSM (Shi et al., 2023) and MSVAMP (Chen et al., 2023) to evaluate the performance of LLMs in multilingual mathematical reasoning. We evaluate in zero-shot chain-of-thought reasoning (Wei et al., 2022) setting. To further assess the task generalization of LinguaLIFT, we incorporate several challenging multilingual datasets, including X-CSQA (Lin et al., 2021) for commonsense reasoning and XNLI (Conneau et al., 2018) for natural language inference.
@@
We utilize English-only instruction data following prior work (Lu et al., 2024; Huang et al., 2024), which include MetaMathQA (Yu et al., 2024) for mathematical reasoning, MultiNLI (Williams et al., 2018) for natural language inference, and a set of unified commonsense reasoning tasks, comprising the X-CSQA, OpenBookQA (Mihaylov et al., 2018), ARC (Clark et al., 2018), and QASC (Khot et al., 2020) datasets. Additionally, to further explore the potential of the proposed method, we incorporate the recent advanced mathematical reasoning instruction dataset, OpenMathInstruct-2 (Toshniwal et al., 2024). Statistics of the datasets involved are presented in Table 5, and the prompts for each task are given in Appendix A.3.
@@ 5.2 Baselines
We consider three categories of baselines: (1) Mono-SFT (Luo et al., 2023; Yue et al., 2024; Yu et al., 2024; Toshniwal et al., 2024; Zhu et al., 2024b), a vanilla method that fine-tunes the model on English task datasets. (2) Multi-SFT (Chen et al., 2023; Zhu et al., 2024b; She et al., 2024), a translation-based method that fine-tunes the model on multilingual task datasets. (3) Leveraging External Tools or Models (Shi et al., 2023; Yoon et al., 2024; Huang et al., 2024), a method that leverages external translation systems or models with existing powerful LLMs. More training details are presented in Appendix A.4.

### v2

# LinguaLIFT: An Effective Two-stage Instruction Tuning Framework for Low-Resource Language Reasoning

@@ Abstract
Large language models (LLMs) have exhibited impressive multilingual reasoning capabilities, driven by extensive multilingual pre-training corpora and instruction fine-tuning data. However, a performance gap exists between high- and low-resource language reasoning tasks due to the language imbalance in the pre-training corpus, which is exacerbated by evaluation bias in existing reasoning benchmarks lacking low-resource language coverage. To alleviate this issue, we propose LinguaLIFT, a two-stage instruction tuning framework for advancing low-resource language reasoning. LinguaLIFT employs a language alignment layer to capture multilingual alignment in a code-switched tuning way without requiring multilingual instruction or parallel data, thereby transferring the cross-lingual reasoning capabilities to low-resource languages through English-only instruction tuning data. To comprehensively evaluate the multilingual reasoning capabilities, we introduce the Multilingual Math World Problem (MMWP) benchmark, which spans 21 low-resource, 17 medium-resource, and 10 high-resource languages. Experimental results show that LinguaLIFT outperforms several competitive baselines across MMWP and four widely used benchmarks.  [^1] Our code and data will be released once accepted.
LinguaLIFT: An Effective Two-stage Instruction Tuning Framework for Low-Resource Language Reasoning
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated impressive reasoning capabilities, opening up a new paradigm for artificial intelligence (Wei et al., 2022; Zhao et al., 2023). These capabilities arise from self-supervised pre-training on large-scale multilingual data, followed by supervised fine-tuning on multilingual instruction data (Ouyang et al., 2022; Muennighoff et al., 2023; Chung et al., 2024). As a result, LLMs have achieved remarkable performance across various natural language processing tasks (Chen et al., 2024a, b; Zhang et al., 2024), especially high-resource language mathematical reasoning tasks (Shi et al., 2023; She et al., 2024).
@@
However, a substantial performance gap still exists between high- and low-resource language reasoning (Shi et al., 2023; Zhu et al., 2024), primarily due to the language imbalance within the large-scale multilingual corpora during pre-training, where high-resource languages dominate (Touvron et al., 2023; Lu et al., 2024). This gap is further exacerbated by evaluation bias inherent in existing multilingual reasoning benchmarks, which exhibit limited coverage of low-resource languages Shi et al. (2023); Chen et al. (2023). As illustrated in Figure  1, while LLMs can answer correctly to the query in English, they struggle with the same query in low-resource languages (e.g., Bengali (Bn), Swahili (Sw)). These discrepancies emphasize LLMs’ fundamental limitation in low-resource language reasoning.
To alleviate this issue, this study proposes a two-stage instruction tuning framework (LinguaLIFT). The first stage introduces a language alignment layer to adapt a pre-trained multilingual encoder, utilizing code-switched translation data to enhance multilingual alignment. Building on this strengthened alignment, the second stage fine-tunes the LLM on English instruction data while keeping the alignment layer frozen, thereby transferring reasoning capabilities from English to low-resource languages. This study builds a new benchmark named Multilingual Math World Problem (MMWP), spanning 21 low-resource, 17 medium-resource, and 10 high-resource languages. Experiments demonstrate that LinguaLIFT significantly outperforms several strong competitive methods on the MMWP and four widely used benchmarks, such as MGSM (Shi et al., 2023), MSVAMP (Chen et al., 2023), XNLI (Conneau et al., 2018) and X-CSQA (Lin et al., 2021).
To summarize, our primary contributions are:
  • The proposed LinguaLIFT significantly enhances low-resource language reasoning without requiring multilingual or parallel data.
  • This study introduces MMWP, a new mathematical reasoning benchmark covering languages at varying resource levels, to evaluate multilingual reasoning tasks comprehensively.
  • Experiments across mathematical and commonsense reasoning demonstrate the effectiveness and generalizability of LinguaLIFT in advancing low-resource language reasoning.
@@ 2 Related Work
Leveraging cross-lingual transfer within multilingual models.
While recent research highlights multilingual models’ robust cross-lingual transfer capabilities (Conneau et al., 2020; FitzGerald et al., 2023; Shaham et al., 2024)—achieved through alignment improvements via multilingual instruction or parallel data (Pfeiffer et al., 2020; Pan et al., 2021; Feng et al., 2022; Li et al., 2023a; Chen et al., 2024c)—such approaches remain impractical in low-resource scenarios due to data scarcity. To address this limitation, LinguaLIFT advances code-switched tuning, demonstrating success in cross-lingual transfer (Yang et al., 2020; Li et al., 2024), to enhance multilingual alignment without requiring multilingual instructions or parallel corpora.
@@
Recent efforts to improve multilingual mathematical reasoning for LLMs can be categorized into three ways: 1) Prompting close-source LLMs: Qin et al. (2023), Shi et al. (2023), and Huang et al. (2023) designed prompts for closed-source LLMs like ChatGPT, translating non-English contexts into English for reasoning. This approach is limited by translation quality and does not improve multilingual understanding or works well for open-source LLMs (Zhu et al., 2024). 2) Instruction-tuning open-source LLMs: MathOctopus (Chen et al., 2023), xCoT (Chai et al., 2024), and mCoT (Lai and Nissim, 2024) adopted a translate-training method, translating English reasoning datasets into non-English and instruction-tuning LLMs. QAlign (Zhu et al., 2024) and MAPO (She et al., 2024) proposed approaches to transfer mathematical reasoning capabilities from English to non-English. While these methods improve multilingual reasoning, they incur high translation costs and errors, making them impractical for low-resource languages. 3) Bridging existing skilled LLMs to multilingualism: Langbridge (Yoon et al., 2024) combines pre-trained multilingual models with skilled reasoning LLMs, but a performance gap persists between low-resource and high-resource languages. While MindMerger (Huang et al., 2024) shows significant improvement, it relies heavily on parallel and multilingual reasoning data.
Figure 2: Overview of the proposed LinguaLIFT approach. Stage-I (Language Align): A language alignment layer is introduced into the LLM to adapt the pre-trained multilingual encoder, thereby enhancing multilingual alignment through code-switched tuning. Stage-II (Task Transfer): The LLM is fine-tuned on high-quality, English-only instruction data with the language alignment layer frozen, allowing the LLM to transfer reasoning capabilities learned from English to low-resource languages.
Unlike prior approaches that depend on costly translated data or impractical parallel corpora in low-resource settings, LinguaLIFT advances low-resource language reasoning without requiring translated multilingual or parallel datasets.
@@ 3 Methodology
To harness cross-lingual transfer for low-resource languages reasoning, this study proposes leveraging the language alignment established by code-switched tuning (Section §3.1) to transfer reasoning capabilities from English to low-resource languages (Section §3.2), as illustrated in Figure 2. Formally, given an input sequence $x$ with $l_{x}$ tokens, a multilingual pre-trained encoder generates a language-agnostic representation $X\in\mathbb{R}^{l_{x}\times d_{1}}$, which is subsequently aligned to the input embedding space of the LLM via a multi-layer perceptron alignment layer, yielding $\hat{X}\in\mathbb{R}^{l_{x}\times d_{2}}$. To leverage the built-in capabilities of the LLM for instruction following, the instruction context $q$ (with $l_{q}$ tokens) is mapped into $Q\in\mathbb{R}^{l_{q}\times d_{2}}$ via the LLM’s embedding layer. The combined input sequence is constructed by concatenating the start token <bos>, $Q$, alignment boundary tokens <enc_start> and <enc_end>, and $\hat{X}$, forming:
@@
This structured input is then processed by the LLM to generate the final response.
3.1 Language Align Stage
Due to the scarcity of parallel corpora in low-resource languages, this study proposes enhancing the model’s multilingual alignment through code-switched tuning. This involves two key parts: (i) multilingual alignment lexicon construction and (ii) code-switched translation instruction tuning.
Multilingual alignment lexicon construction.
An unsupervised word translation method MUSE (Lample et al., 2018) is adopted to construct the multilingual alignment lexicons without parallel data. Specifically, a set of unique words $\mathcal{W}=\{w\}_{i}^{N}$ is extracted from English monolingual corpus $\mathcal{D}$, where $N$ is the number of unique words, excluding named entities and stop words. Word translations are obtained by retrieving the top nearest neighbor via projection of word embeddings into the target space. Further implementation details and validation steps in Appendix A.1 ensure the robustness and reliability of our constructed lexicons.
Code-switched translation instruction tuning.
The constructed multilingual alignment lexicons are used to generate code-switched inputs for instruction tuning. Formally, our training objective is to predict the English translation response tokens $y$ given the instruction context $q$ and code-switched input tokens $x_{l}$ in language $l$. The input embeddings for this task, represented as $(Q,\hat{X_{l}})$ from Equation (1), serve as the LLMs’ input. The optimization objective is defined as follows:
@@ 3.1 Language Align Stage
where $T$ is the target sequence length, $y_{<i}$ represents the preceding tokens, $y_{i}$ denotes the $i$-th token, and $L$ is the set of target languages. Trainable parameters include $\theta$ (language alignment layer and boundary tokens), while $\psi$ (multilingual encoder) and $\phi$ (LLM) remain frozen.
3.2 Task Transfer Stage
Following the language alignment established in the Language Align Stage, the LLM is fine-tuned using high-quality English reasoning instruction data while keeping the language alignment layer frozen. This strategy preserves the acquired multilingual alignment and cross-lingual transfer capabilities, unlocking the potential of LLM to transfer strong reasoning capabilities learned from English to low-resource languages. The training minimizes the negative log-likelihood of the response $y$, given the combined instruction context and query input $(Q,\hat{X}_{en})$:
$-\underset{\phi}{\arg\min}\sum_{i=1}^{T}\log p_{\tilde{\theta},\psi,\phi}\left(y_{i}\mid(Q,\hat{X}_{en}),y_{<i}\right),$ (3)
where $\phi$ denotes the trainable LLM parameter, and $\tilde{\theta}$ is initialized from the checkpoint of $\theta$ trained at the first stage. The multilingual encoder parameters $\psi$ and $\tilde{\theta}$ are frozen throughout the training.
@@ 4 The MMWP Benchmark
Our study employs a typologically diverse sample of 48 languages representing 12 language families and 12 writing systems, encompassing low-, medium-, and high-resource levels. The selection comprises 21 low-resource, 17 medium-resource, and 10 high-resource languages. Language categorization list and selection criteria are detailed in Appendices B.1 and B.2, respectively.
Translation and post-calibration.
English questions are translated into target languages using Google Translate. Translation quality is verified through combined automated and human evaluations. Five annotators then refine inaccuracies via GPT-4-guided post-calibration, followed by reviews and adjustments from two language experts prominent in minor languages. Detailed accuracy assurance procedures are provided in Appendix B.3.
@@ 5.1 Datasets
We use the MMWP and the latest multilingual benchmarks, MGSM (Shi et al., 2023) and MSVAMP (Chen et al., 2023) to evaluate the performance of LLMs in multilingual mathematical reasoning through zero-shot chain-of-thought reasoning (Wei et al., 2022) setting. To further assess the task generalization of LinguaLIFT, we incorporate several challenging multilingual datasets, including X-CSQA (Lin et al., 2021) for commonsense reasoning and XNLI (Conneau et al., 2018) for natural language inference.
@@
We utilize English-only instruction data following prior work (Lu et al., 2024; Huang et al., 2024), which include MetaMathQA (Yu et al., 2024) for mathematical reasoning, MultiNLI (Williams et al., 2018) for natural language inference, and a set of unified commonsense reasoning tasks, comprising the X-CSQA, OpenBookQA (Mihaylov et al., 2018), ARC (Clark et al., 2018), and QASC (Khot et al., 2020) datasets. Additionally, to further explore the potential of the proposed method, we incorporate the recent advanced mathematical reasoning instruction dataset, OpenMathInstruct-2 (Toshniwal et al., 2024). Statistics of the datasets involved are presented in Appendix A.2, and the prompts for each task are given in Appendix A.3.
@@ 5.2 Baselines
Three categories of baselines are considered: (1) Mono-SFT: MAmmoTH (Yue et al., 2024), WizardMath (Luo et al., 2023), MetaMath (Yu et al., 2024), and OpenMath2 (Toshniwal et al., 2024) fine-tune on respective English instruction dataset. (2) Multi-SFT: MathOctopus (Chen et al., 2023), QAlign (Zhu et al., 2024), MAPO (She et al., 2024) fine-tune on translated multilingual instruction datasets (e.g., MGSM8K-Instruct (Chen et al., 2023)) or parallel corpus. (3) Leveraging External Tools or Models: Translate-En (Shi et al., 2023) use SOTA translation models (Team et al., 2022) to translate the query into English; LangBridge (Yoon et al., 2024) and MindMerger (Huang et al., 2024) replace the input of LLMs with the hidden states output by mT5 (Xue et al., 2021). To ensure a fair comparison with MindMerger (the previous SOTA), we further consider continuing training (Add Low-Resource) or retraining (Low-Resource Retrain) with open-source low-resource parallel data. More training details are presented in Appendix A.4.